uiek ratio to exceed the current ratio? iy moou the current ratio? 15. Interpret a working capital requirement of S0. PROBLEMS For the problems, CCC-cash conversion cycle, DSo days sales outstanding, DIH ld, and DPO days payable outstanding. 1.) Suppose that a firm has a 30-day DSO. Determine the firms DIlH if the operating cycle is days inventory a. 30 days b. 60 days c. 80 days 2. Suppose that a firm has a 50-day DIH and a 30-day DPO. Determine the firms DSO if the CCc is: a. 90 days b. 60 days c. 20 days Suppose that a firm has a 60-day operating cycle. Determine the firms DPO if the CCCis: a. 40 days b. 0 days c. 20 days 3. COGS-45% of revenues. Receivables-$1,100,000, payable: $700,000 and inventory is $800,000. a. Calculate the operating cycle. b. Calculate the CCC c. With all else constant, what level of payables will llow the firm to have a CCC of o days? d. With all else constant, what level of inventory will llow the firm to have a CCC of 0 days? e. With all else constant, what level of receivables will allow the firm to have a CCC of 0 days? 4. Use the following financial values to answer the questions below. Revenues $4,200,000 and Is your answer feasible? page no
Answer #1

Answer 1:

a. DIH = 0 days

b. DIH = 30 days

c. DIH = 50 days

Workings:

DIH = Operating cycle - DSO

a. DIH = 30 - 30 = 0 days

b. DIH = 60 - 30 = 30

C. DIH =80- 30 = 50

Answer 2:

a. DSO = 70 days

b. DSO = 40 days

c. DSO = 0 days

Workings:

DSO = CCC - DIH + DPO

a. DSO = 90 - 50 + 30 = 70

b. DSO = 60- 50 + 30 = 40

c. DSO = 20- 50 + 30 = 0

Answer 3:

a. DPO = 20 days

b. DPO = 60 days

c. DPO = 40 days

Workings:

DPO = Operating cycle - CCC

a. DPO = 60 - 40 = 20 days

b. DPO = 60 - 0 = 60

C. DPO = 60- 20 = 40

Answer 4:

DSO = Receivable / (sales / 365) = 1100000 / (4200000 / 365) = 96 days

DIH = Inventory/ (COGS /365) = 800000 /(45% * 4200000 / 365) = 154 days

DPO = 700000 / (45% * 4200000 / 365) = 135 days

a. Operating cycle = DSO +DIH = 96 + 154 = 250 days

b. CCC = Operating cycle - DPO = 250 - 135 = 115 days

c. CCC will be 0 days if DPO = Operating cycle = 250 days

Payables = 250 * (45% * 4200000 / 365) =$1,294,521

d. CCC will be 0 when:

DIH = CCC - DSO + DPO = 0 - 96 + 135 =39 days

Inventory =39 * (45% * 4200000 / 365) = $201,945

e. CCC will be 0 when:

DSO = CCC - DIH + DPO = 0 - 154 + 135 = -19 days

DSO has to be negative to have CCC = 0 (all else remaining same). This is not feasible.